Calculate the total mass, in grams of precipitate produced when 10.0 ml of 2.50 M SrS (aq) reacts completely with ZnSO4 (aq).


1
Expert's answer
2021-04-25T23:10:03-0400

SrS + ZnSO4 --> ZnS + SrSO4

Molar Mass of SrS = 119.68

= 119.68 × 2.5/10 = 29.92g

Molar Mass of ZnSO4 = 161.47g

= 161.47 × 2.5/10 = 40.3675g

= 29.92 + 40.3675

= 70.2875g